A steel lift column in a service station is 3.0 meters long and. has an area of 0.2 m2. Young’s modulus for steel is 20 x 1010 N/m2. By how much does the column shrink when a 3000 kg load is on it?

A)3.1 x 10-5 m

B)2.25 x 10-7 m

C)3.5 x 10-6 m

D)2.2 x 10-6 m
Answer #1

1.
σ = 3000*9.8/(π*.10²) = .935E6 Pa
ε = σ/Y = 4.675E-6
dL = ε*L = 1.4025E-5 m